OpenStudy (anonymous):

which calues of x and y will make the expression 3(-2x-y)^2 equal to 3?

OpenStudy (anonymous):

i mean values

OpenStudy (ranga):

when (-2x-y) = 1 or -1

OpenStudy (anonymous):

-1

OpenStudy (ranga):

Since (-2x-y) is squared in the expression, it can be -1 or +1 and the expression will equal 3.

OpenStudy (ranga):

3(-2x-y)^2 = 3 divide both sides by 3 (-2x-y)^2 = 1 take square root on both sides (-2x-y) = sqrt(1) = -1 or +1 -2x - y = -1 or -2x - y = 1